40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 13 03:00:01 2011, Jenne, Update, IOO, WFS shenanigans D990249-B-1_MCWFS_schematic.pdf
    Reply  Thu Jan 13 12:56:57 2011, rana, Update, IOO, WFS shenanigans PerkinElmerQPDs.pdf
    Reply  Fri Jan 28 13:07:31 2011, Jenne, Update, IOO, Beam is back on the WFS 
    Reply  Fri Jan 28 18:19:21 2011, Jenne, Update, IOO, WFS2 has some kind of oil on it Oily_WFS2.jpg
    Reply  Mon Feb 14 15:59:49 2011, Jenne, Update, IOO, WFS quantum efficiency as a function of angle 
       Reply  Wed Jun 15 00:49:34 2011, Suresh, Update, IOO, WFS2 has been fixed. WFS_QE_measurement.pngSensorsBrochure-p12.pdfP6150121.JPGP6150124.JPG
Message ID: 4222     Entry time: Fri Jan 28 13:07:31 2011     In reply to: 4148
Author: Jenne 
Type: Update 
Category: IOO 
Subject: Beam is back on the WFS 

The MC WFS have had beam dumps in front of them for the past ~2 weeks, until I could find the appropriate optic to put in the WFS path, to avoid melting the WFS' electronics. 

Koji noted that Steve had a W2-45S in a secret stash near his desk (which Steve later had put into the regular optics storage shelves down the Yarm), so I used that in front of the black hole beam dump on the AS table.  Now the beam is ~1W reflected from the unlocked mode cleaner, and ~100mW goes to the MC REFL PD.  The other 900mW now goes to this W2, and only ~5mW is reflected toward the MC WFS.  Most of the 900mW is transmitted through the window and dumped in the black hole.  There is a ghost beam which is reflected off the back surface of the wedged window, and I have blocked this beam using a black anodized aluminum dump.  I will likely change this to a razor dump if space on the table allows.  I have aligned the beam onto WFS1 and WFS2, although I did not re-align the mode cleaner first, so this alignment of the WFS will likely need to be redone. 

WFS1 has about 2mW incident, and WFS2 has about 3mW incident, when the mode cleaner is unlocked.  I have not yet measured the power incident when the MC is locked, although obviously it will be much smaller.

Except that I might temporarily remove one of the WFS for more quantum efficiency measurements later today, the WFS should be ready to turn back on for alignment stabilization of the mode cleaner. 

Quote:

My goal this afternoon was to measure the quantum efficiency of the MC WFS.  In the process of doing this, I discovered that when I reverted a change in the MCWFS path (see elog 4107 re: this change), I had not checked the max power going to the WFS when the MC unlocks.

Current status:

MC locks (is locked now).  No light going to WFS at all (to prevent MC WFS french-fry action).  Quantum Efficiency measured.

The Full Story:

Power to WFS:

Rana asked me to check out the quantum efficiency of the WFS, so that we can consider using them for aLIGO.  This involves measuring the power incident on the PDs, and while doing so, I noticed that WFS1 had ~160mW incident and WFS2 had ~240mW incident while the mode cleaner was unlocked.  This is bad, since they should have a max of ~10mW ever.  Not that 200mW is going to destroy the PD immediately, but rather the current out, with the 100V bias that the WFS have, is a truckload of power, and the WFS were in fact getting pretty warm to the touch.  Not so good, if things start melting / failing due to extended exposure to too much heat.

The reason so much power was going to the WFS is that it looks like Yuta/Koji et. al., when trying to use the WFS as a MC1 oplev, changed out 2 of the beam splitters in the MC WFS / MC Refl path, not just one.  Or, we've just been crispy-frying our WFS for a long time.  Who knows?  If it is option A, then it wasn't elogged.  The elog 3878 re: BS changeout only mentions the change of one BS.

Since the MC Refl path has a little more than ~1W of power when the MC is unlocked, and the first BS (which was reverted in elog 4107) is a 10% reflector, so ~100mW goes to the MC Refl PD, and ~900mW goes to the MC WFS path.  In front of a Black Hole beam dump was sitting a BS1-33, so we were getting ~300mW reflected to be split between the 2 WFS, and ~600mW dumped.  The new plan is to put a W2 window in place of this BS1-33, so that we get hopefully something like 0.1% reflected toward the WFS, and everything else will be dumped.  I could not find a W2-45S (everything else is S, so this needs to be S as well).  I found a bunch of W2-0deg, and a few W2-45P.  Does anyone have a secret stash of W2-45S's???  To avoid any more excessive heat just in case, for tonight, I have just left out this mirror entirely, so the whole MC WFS beam is dumped in the Black Hole.  The WFS also have aluminum beam dumps in front of them to prevent light going in.  None of this affects the MC Refl path, so the MC can still lock nice and happily.

 

ELOG V3.1.3-